Find the solution of 1/5 + 3/10 algorithmically and diagrammatically
1
Expert's answer
2020-10-26T20:29:31-0400

"\dfrac{1}{5} + \dfrac{3}{10} ="


"[\dfrac{1}{5} \times \dfrac{2}{2}] + [\dfrac{3}{10}] ="


"\dfrac{2}{10} + \dfrac{3}{10} ="


"\dfrac{5}{10} = \dfrac{1}{2}"